# Deformable One-shot Face Stylization via DINO Semantic Guidance

### 01 Mar 2024 v1

@@ Abstract
Figure 1: We propose a deformation-aware face stylization framework trained on a single real-style image pair (left). Our framework can generate diverse, high-quality, stylized faces with desired deformations, while maintaining the input identity well (right).
footnotetext: *Corresponding author
@@ 1 Introduction
Can we stylize facial images according to only one style example while simultaneously considering appearance change and structure exaggeration? Previous methods [6, 45, 21] solve this problem by building “fake” guidance across the source and target domain based on the inversion of the style reference. However, as shown in Fig. 2, current GAN inversion techniques, such as [2] and [32], still can not produce faithful mappings for artistic images, which will definitely mislead the geometric deformation, especially when strong exaggerations occur. We argue that we can build reliable deformation guidance across domains using a real-style image pair instead of only the style example, thus reducing the difficulty; see, e.g., Fig. 1. Nonetheless, existing one-shot methods still failed to capture the deformation pattern because of the weaker structural guidance they employed; see Sec. 5 for more details.
@@ 4.2 DINO-based domain adaption
As a few-shot training, adapting the generator only with the above directional deformation guidance will easily result in over-fitting. Inspired by [24], we introduce a relative structural consistency to preserve the structural diversity across domains. Specifically, we measure the difference/similarity between every two generated batch samples of the same domain. The self-similarity vectorThe self-similarity between DINO tokens forms a large similarity matrix (in size of 783*783). We flatten it into a self-similarity vector. computed with DINO tokens is used as the structure representation for each sample, which is $s_{i}=\frac{E_{D}(I_{i})E_{D}(I_{i})^{T}}{||E_{D}(I_{i})||^{2}_{2}}$. For the ground truth pair, we involve them in the relative structural consistency as well. Accordingly, two relative similarity matrices can be formed for the source and target domain, respectively, as depicted in Fig. 6. We then transform them into probability distributions by
@@ Acknowledgements
This work was supported in parts by NSFC (U21B2023, U2001206, 62161146005), GD Natural Science Foundation (2022A1515010221), DEGP Innovation Team Program (2022KCXTD025), Shenzhen Science and Technology Program (KQTD20210811090044003, RCJC20200714114435012), and Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ).

### 04 Mar 2024 v2

@@ 1 Introduction
Can we stylize facial images according to only one style example while simultaneously considering appearance change and structure exaggeration? Previous methods [6, 45, 21] solve this problem by building “fake” guidance across the source and target domain based on the inversion of the style reference. However, as shown in Fig. 1, current GAN inversion techniques, such as [2] and [32], still can not produce faithful mappings for artistic images, which will definitely mislead the geometric deformation, especially when strong exaggerations occur. We argue that we can build reliable deformation guidance across domains using a real-style image pair instead of only the style example, thus reducing the difficulty; see, e.g., Fig. LABEL:fig:teaser. Nonetheless, existing one-shot methods still failed to capture the deformation pattern because of the weaker structural guidance they employed; see Sec. 5 for more details.
@@ 4.2 DINO-based domain adaption
As a few-shot training, adapting the generator only with the above directional deformation guidance will easily result in over-fitting. Inspired by [24], we introduce a relative structural consistency to preserve the structural diversity across domains. Specifically, we measure the difference/similarity between every two generated batch samples of the same domain. The self-similarity vector***The self-similarity between DINO tokens forms a large similarity matrix (in size of 783*783). We flatten it into a self-similarity vector. computed with DINO tokens is used as the structure representation for each sample, which is $s_{i}=\frac{E_{D}(I_{i})E_{D}(I_{i})^{T}}{||E_{D}(I_{i})||^{2}_{2}}$. For the ground truth pair, we involve them in the relative structural consistency as well. Accordingly, two relative similarity matrices can be formed for the source and target domain, respectively, as depicted in Fig. 5. We then transform them into probability distributions by
@@ Acknowledgements
This work was supported in parts by NSFC (U21B2023, U2001206, 62161146005), GD Natural Science Foundation (2022A1515010221), DEGP Innovation Team Program (2022KCXTD025), Shenzhen Science and Technology Program (KQTD20210811090044003, RCJC20200714114435012), Guangdong Laboratory of Artificial Intelligence and Digital Economy (SZ) and Scientific Development Funds of Shenzhen University.